OpenStudy (anonymous):

Factor by grouping. x3 + 5x2 - 2x - 10

OpenStudy (bahrom7893):

\[x^3+5x^2-2x-10=x^2(x+5)-2(x+5)=(x+5)(x^2-2)=(x+5)(x-\sqrt(2)(x+\sqrt(2)\]

OpenStudy (bahrom7893):

\[(x+5)(x-\sqrt(2)(x+\sqrt(2)\]

OpenStudy (anonymous):

\[x^2(x+5)-2(x+5)=(x^2-2)(x+5)\]